# Faster Convergence of Riemannian Stochastic Gradient Descent with Increasing Batch Size

### v3

@@ 3 Convergence Analysis
First, we prepare some lemmas for the convergence analysis of RSGD.
Lemma 1 (Descent Lemma).
Let $(x_{t})_{t}$ be a sequence generated by RSGD and $(\eta_{t})_{t}$ be a positive valued sequence. Then, under Assumptions 1 and 2, we obtain
Lemma 1 implies Lemma 2 which plays a central role over the following convergence analyses. The proofs of both lemmas are given in Appendices C.1 and C.2.
Lemma 2 (Underlying Analysis).
Let $(x_{t})_{t}$ be a sequence generated by RSGD and $\eta_{\max}>0$. We consider a positive valued sequence $(\eta_{t})_{t}$ such that $\eta_{t}\in[0,\eta_{\max}]\subset[0,\frac{2}{L_{r}})$. Then, under Assumptions 1 and 2, we obtain
@@ 3.1 Case (i): Constant BS; Constant or Decaying LR
In this case, we consider a BS $(b_{t})_{t}$ and a LR $(\eta_{t})_{t}$ such that $b_{t}=b$ and $\eta_{t+1}\leq\eta_{t}$. In particular, we set the following examples with constant or decaying LRs.
$\text{Constant LR:}\quad\eta_{t}=\eta_{\max},$ (1)
$\text{Diminishing LR:}\quad\eta_{t}=\frac{\eta_{\max}}{\sqrt{t+1}},$ (2)
$\text{Cosine Annealing LR:}\quad\eta_{t}\coloneq\eta_{\min}+\frac{\eta_{\max}-\eta_{\min}}{2}\left(1+\cos\frac{t}{T}\pi\right),$ (3)
$\text{Polynomial Decay LR:}\quad\eta_{t}\coloneq\eta_{\min}+(\eta_{\max}-\eta_{\min})\left(1-\frac{t}{T}\right)^{p},$ (4)
where $\eta_{\max}$ and $\eta_{\min}$ are positive values satisfying $0\leq\eta_{\min}\leq\eta_{\max}<\frac{2}{Lr}$. Note that $\eta_{\max}$ (resp. $\eta_{\min}$) becomes the maximum (resp. minimum) of $\eta_{t}$; namely, $\eta_{t}\in[\eta_{\min},\eta_{\max}]\subset[0,\frac{2}{L_{r}})$.
Theorem 1.
We consider LRs equation 1, equation 2, equation 3, and equation 4 and a constant BS $b_{t}=b>0$ under the assumptions of Lemma 2. Then, we obtain
  • Diminishing LR equation 2
    $\min_{t\in\{0,\cdots,T-1\}}\mathbb{E}[\|\mathrm{grad}f(x_{t})\|_{x_{t}}^{2}]$ $\leq\frac{Q_{1}+Q_{2}\sigma^{2}b^{-1}\log T}{\sqrt{T}},$
  • Otherwise equation 1, equation 3, equation 4
    $\min_{t\in\{0,\cdots,T-1\}}\mathbb{E}[\|\mathrm{grad}f(x_{t})\|_{x_{t}}^{2}]$ $\leq\frac{\tilde{Q}_{1}}{T}+\frac{\tilde{Q}_{2}\sigma^{2}}{b},$
where $Q_{1},Q_{2},\tilde{Q}_{1}$, and $\tilde{Q}_{2}$ are constants which do not depend on $T$.
Therefore, a diminishing LR equation 2 has $\|G_{T}\|=O\left(\frac{\sqrt{\log T}}{\sqrt[4]{T}}\right)$, whereas equation 1, equation 3, and equation 4 have $\|G_{T}\|=O\left(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}}\right)$. The proof is in Appendix C.3.
@@ 3.2 Case (ii): Increasing BS; Constant or Decaying LR
where $\gamma,c>1$, and $a>0$. One can easily check this. (For example, $\frac{m}{m+1}\leq\left\lceil\frac{t}{(m+1)K}\right\rceil<1$ $(\forall t\in S_{m})$ holds.)
Theorem 2.
We consider BSs equation 5 and equation 6 and LRs equation 1, equation 2, equation 3, and equation 4 under the assumptions of Lemma 2. Then, the following holds for both constant and increasing BSs.
  • Diminishing LR equation 2
    $\min_{t\in\{0,\cdots,T-1\}}\mathbb{E}[\|\mathrm{grad}f(x_{t})\|_{x_{t}}^{2}]$ $\leq\frac{Q_{1}+Q_{2}\sigma^{2}b_{0}^{-1}}{\sqrt{T}},$
  • Otherwise equation 1, equation 3, equation 4
    $\min_{t\in\{0,\cdots,T-1\}}\mathbb{E}[\|\mathrm{grad}f(x_{t})\|_{x_{t}}^{2}]$ $\leq\frac{\tilde{Q}_{1}+\tilde{Q}_{2}\sigma^{2}b_{0}^{-1}}{T},$
where $Q_{1},Q_{2},\tilde{Q}_{1}$, and $\tilde{Q}_{2}$ are constants which do not depend on $T$.
Therefore, the diminishing LR equation 2 has $\|G_{T}\|=O(\frac{1}{\sqrt[4]{T}})$, and the other LRs equation 1, equation 3, and equation 4 have $\|G_{T}\|=O(\frac{1}{\sqrt{T}})$. The proof is in Appendix C.4.
@@ 3.3 Case (iii): Increasing BS; Warm-up Decaying LR
In this case, we consider a BS $(b_{t})_{t}$ and a LR $(\eta_{t})_{t}$ such that $b_{t}\leq b_{t+1}$ and $\eta_{t}\leq\eta_{t+1}$ $(t\leq T_{w}-1),\eta_{t+1}\leq\eta_{t}$ $(t\geq T_{w})$. As examples of an increasing warm-up LR, we use an exponentially increasing LR and a polynomially increasing LR, both increasing every $K^{\prime}$ steps. We set $K$, which was defined in Case (ii) (Section 3.2), as $lK^{\prime}$, where $l\in\mathbb{N}$ (thus, $K>K^{\prime}$). Namely, we consider a setting in which the BS is increased every $l$ times the LR is increased. To formulate examples of an increasing LR, we define $M^{\prime}\coloneq\lfloor\frac{T}{K^{\prime}}\rfloor$ and formalize the LR: for every $m\in\{0,\ldots,M^{\prime}-1\},t\in S^{\prime}_{m}\coloneq[mK^{\prime},(m+1)K^{\prime})\cap\mathbb{N}$ $(S^{\prime}_{0}\coloneq[0,K^{\prime}))$,
$\text{Exponential Growth LR:}\quad\eta_{t}\coloneq\eta_{0}\delta^{m\left\lceil\frac{t}{(m+1)K^{\prime}}\right\rceil},$ (7)
$\text{Polynomial Growth LR:}\quad\eta_{t}\coloneq\left(sm\left\lceil\frac{t}{(m+1)K^{\prime}}\right\rceil+\eta_{0}\right)^{q},$ (8)
where $s>0$ and $q>1$. Furthermore, we choose $\gamma,\delta>1$, and $l\in\mathbb{N}$ such that $\delta^{2l}<\gamma$ holds. Additionally, we set $l_{w}\in\mathbb{N}$ such that $T\geq T_{w}\coloneq l_{w}K^{\prime}\geq lK^{\prime}$. The examples of an increasing BS used in this case are an [exponential growth BS] equation 5 and a polynomial growth BS equation 6. As examples of a warm-up LR, we consider LRs that are increased using exponential growth LR equation 7 and polynomial growth LR equation 8 corresponding to the exponential and polynomial growth BSs for the first $T_{w}$ steps and then decreased using a constant LR equation 1, a diminishing LR equation 2, a cosine annealing LR equation 3, and a polynomial decay LR equation 4 for the remaining $T-T_{w}$ steps. Note that $\eta_{\max}\coloneq\eta_{T_{w}-1}$.
Theorem 3.
We consider BSs equation 5 and equation 6 and warm-up LRs equation 7 and equation 8 with decay parts equation 1, equation 2, equation 3, and equation 4 under the assumptions of Lemma 2. Then, the following holds for both constant and increasing BSs.
  • Decay part: Diminishing equation 2
  • Decay part: Otherwise equation 1, equation 3, equation 4
where $Q_{1},Q_{2},\tilde{Q}_{1}$, and $\tilde{Q}_{2}$ are constants which do not depend on $T$.
Therefore, the diminishing LR equation 2 has $\|G_{T}\|=O(\frac{1}{\sqrt{\sqrt{T+1}-\sqrt{T_{w}+1}}})$, and the other LRs equation 1, equation 3, and equation 4 have $\|G_{T}\|=O(\frac{1}{\sqrt{T-T_{w}}})$. The proof is in Appendix C.5.
@@ 3.4 Case (iv): Constant BS; Warm-up Decaying LR
In this case, we consider a BS $(b_{t})_{t}$ and a LR $(\eta_{t})_{t}$ such that $b_{t}=b$ and $\eta_{t}\leq\eta_{t+1}$ $(t\leq T_{w}-1),\eta_{t+1}\leq\eta_{t}$ $(t\geq T_{w})$. As examples of a warm-up LR, we use an exponential growth LR equation 7 and a polynomial growth LR equation 8 for the first $T_{w}$ steps and then a constant LR equation 1, diminishing LR equation 2, cosine annealing LR equation 3, or polynomial decay LR equation 4 for the remaining $T-T_{w}$ steps. The other conditions are the same as in Case (iii) (Section 3.3).
Theorem 4.
We consider a constant BS $b_{t}=b>0$ and warm-up LRs equation 7 and equation 8 with decay parts equation 1, equation 2, equation 3, or equation 4 under the assumptions of Lemma 2. Then, we obtain
  • Decay part: Diminishing equation 2
  • Decay part: Otherwise equation 1, equation 3, equation 4
where $Q_{1},Q_{2},\tilde{Q}_{1}$, and $\tilde{Q}_{2}$ are constants which do not depend on $T$.
Therefore, the diminishing LR equation 2 has $\|G_{T}\|=O\left(\sqrt{\frac{\log T-\log T_{w}}{\sqrt{T+1}-\sqrt{T_{w}+1}}}\right)$, and the other LRs equation 1, equation 3, and equation 4 have $\|G_{T}\|=O\left(\sqrt{\frac{1}{T-T_{w}}+\frac{\sigma^{2}}{b}}\right)$. The proof is in Appendix C.6.
@@ 4 Numerical Experiment
We experimentally evaluated the performance of RSGD for the two types of BSs and various types of LRs introduced in Section 3. They were run on an iMac (Intel Core i5, 2017) running the macOS Venture operating system (ver. 13.7.1). The algorithms were written in Python (3.12.7) using the NumPy (1.26.0) and Matplotlib (3.9.1) packages. The Python codes are available at https://anonymous.4open.science/r/faster_rsgd-CE24. We set $p=2.0$ in equation 4 and $\eta_{\min}:=0$. In Cases (i) and (ii), we used an initial LR $\eta_{\max}$ selected from $\{0.5,0.1,0.05,0.01,0.005\}$. In Case (ii), we set $K=1000,\gamma=3.0$ and $a=c=2.0$. The plots of the objective function values corresponding to this section are in Appendix B and Appendix D. The results for Cases (iii) and (iv) are given in Appendix E.
@@ 4.1 Principal Component Analysis
Figure 1: Norm of the gradient of the objective function against number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 on COIL100 dataset.
Figure 2: Norm of the gradient of the objective function against number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 on MNIST dataset.
We can formulate the principal component analysis (PCA) problem as an optimization problem on the Stiefel manifold Kasai et al. (2019); for a given dataset $\{x_{j}\}_{j=1,\ldots,N}\subset\mathbb{R}^{n}$ and $r$ $(\leq n)$,
minimize $f(U)\coloneq\frac{1}{N}\sum_{j=1}^{N}\|x_{j}-UU^{\top}x_{j}\|^{2},$
subject to $U\in\mathrm{St}(r,n)\coloneq\{U\in\mathbb{R}^{r\times n}\mid U^{\top}U=I_{n}\}.$
We set $r=10$ and used the COIL100 Nene et al. (1996)) and MNIST LeCun et al. (1998) datasets. The Columbia Object Image Library (COIL100) dataset contains $7200$ color camera images of $100$ objects ($72$ poses per object) taken from different angles. We resized the images to $32\times 32$ pixels and transformed each one into a $1024(=32^{2})$ dimensional vector. Hence, we set $(N,n,r)=(7200,1024,10)$. The MNIST dataset contains $60,000$ $28\times 28$-pixel gray-scale images of handwritten numbers 0 to $9$. We transformed each image into a $784(=28^{2})$ dimensional vector and normalized each pixel to the range $[0,1]$. Hence, we set $(N,n,r)=(60000,784,10)$. Furthermore, we used a constant BS with $b_{t}\coloneq 2^{10}$, an exponential growth BS with an initial value $b_{0}\coloneq 3^{5}$, and a polynomial growth BS with an initial value $b_{0}\coloneq 30$.
@@ 4.2 Low-rank Matrix Completion
Figure 3: Norm of the gradient of the objective function against number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 on MovieLens-1M dataset.
@@
Figure 4: Norm of the gradient of the objective function against number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 on Jester dataset.
The LRMC problem involves completing an incomplete matrix $Z=(z_{1},\ldots,z_{N})\in\mathbb{R}^{n\times N}$; $\Omega$ denotes a set of indices for which we know the entries in $Z$. For $a\in\mathbb{R}^{n}$, we define $P_{\Omega_{i}}(a)$ such that the $j$-th element is $a_{j}$ if $(i,j)\in\Omega$ and 0 otherwise. For $U\in\mathbb{R}^{n\times r},z\in\mathbb{R}^{n}$, $q_{j}(U,z):=\mathrm{argmin}_{a\in\mathbb{R}^{r}}\|P_{\Omega_{j}}(Ua-z)\|$. We can now formulate the LRMC problem as the following optimization problem on the Grassmann manifold Boumal & Absil (2015);
minimize $f(U)\coloneq\frac{1}{N}\sum_{j=1}^{N}\|P_{\Omega_{j}}(Uq_{j}(U,z_{j})-z_{j})\|^{2},$
subject to $U\in\mathrm{Gr}(r,n)\coloneq\text{St}(r,n)/\text{O}(r),$
@@
where $\text{O}(r)$ is the $r$-dimensional orthogonal group. We set $r=10$ and used the MovieLens-1M Harper & Konstan (2016) and Jester datasets Goldberg et al. (2001). The MovieLens-1M dataset contains $1,000,209$ ratings given by $6040$ users on $3952$ movies. Every rating lies in $[0,5]$. We normalized each rating to the range $[0,1]$. Hence, we set $(N,n,r)=(3952,6040,10)$. The Jester dataset contains ratings of $100$ jokes from $24,983$ users. Every rating lies in the range [10,10][-10,10]. Hence, we set $(N,n,r)=(24983,100,10)$. Furthermore, we used a constant BS with $b_{t}\coloneq 2^{8}$, an exponential growth BS with an initial value $b_{0}\coloneq 3^{4}$, and an polynomial growth BS with an initial value $b_{0}\coloneq 14$.
@@
The performance in terms of the gradient norm of the objective function against the number of iterations for LRs equation 1, equation 2, equation 3, and equation 4 on the COIL100, MNIST, MovieLens-1M, and Jester datasets are plotted in Figures 1, 2, 3, and 4, respectively. The performance in terms of achieving a small gradient norm was better with an increasing BS than with a constant BS. Even in increasing BSs, an exponential growth BS showed better performance than a polynomial BS. Since the magnitude of increase in an exponential BS is $O(\gamma^{m})$ and it of a polynomial BS is $O(m)$, this numerical results demonstrated that a larger increase in BS leads to better performance. Although the differences in the objective function values are small (one possible reason for this is that the objective function may be flat around the optimal solution), the performance of an increasing BS was equal to or better than a constant BS.
@@
4.3 Comparison of Computational Time between Constant BS and Increasing BS via SFO Complexity

### v4

@@ 3 Convergence Analysis
We begin by presenting a lemma that will be used in the convergence analysis of RSGD. Lemma 3.1 plays a central role is the following convergence analyses; its proof is provided in Appendix D.1.
Lemma 3.1 (Underlying Analysis).
Let $(x_{t})_{t}$ be a sequence generated by RSGD and let $\eta_{\max}>0$. Consider a positive-valued sequence $(\eta_{t})_{t}$ such that $\eta_{t}\in[0,\eta_{\max}]\subset[0,\frac{2}{L_{r}})$. Then, under Assumptions 2.1 and 2, we obtain
@@ 3.1 Case (i): Constant BS; Constant or Decaying LR
In this case, we consider a BS $(b_{t})_{t}$ and an LR $(\eta_{t})_{t}$ such that $b_{t}=b$ and $\eta_{t+1}\leq\eta_{t}$. In particular, we present the following examples with constant or decaying LRs.
$\text{Constant LR:}\qquad\eta_{t}=\eta_{\max},$ (1)
$\text{Diminishing LR:}\qquad\eta_{t}=\frac{\eta_{\max}}{\sqrt{t+1}},$ (2)
$\text{Cosine Annealing LR:}\qquad\eta_{t}\coloneq\eta_{\min}+\frac{\eta_{\max}-\eta_{\min}}{2}\left(1+\cos\frac{t}{T}\pi\right),$ (3)
$\text{Polynomial Decay LR:}\qquad\eta_{t}\coloneq\eta_{\min}+(\eta_{\max}-\eta_{\min})\left(1-\frac{t}{T}\right)^{p},$ (4)
where $\eta_{\max}$ and $\eta_{\min}$ are positive values satisfying $0\leq\eta_{\min}\leq\eta_{\max}<\frac{2}{Lr}$. Note that $\eta_{\max}$ (resp. $\eta_{\min}$) becomes the maximum (resp. minimum) value of $\eta_{t}$; namely, $\eta_{t}\in[\eta_{\min},\eta_{\max}]\subset[0,\frac{2}{L_{r}})$.
Theorem 3.2.
We consider LRs (1), (2), (3), and (4) and a constant BS $b_{t}=b>0$ under the assumptions of Lemma 3.1. Then, we obtain
[…]
where $Q_{1},Q_{2},\tilde{Q}_{1}$, and $\tilde{Q}_{2}$ are constants that do not depend on $T$.
@@ 3.2 Case (ii): Increasing BS; Constant or Decaying LR
where $\gamma,c>1,a>0$, and $\mathbb{N}_{0}\coloneq\mathbb{N}\cup\{0\}$.
Theorem $3.$
We consider BSs (5) and (6) together with LRs (1), (2), (3) and (4) under the assumptions stated in Lemma 3.1. Then, the following results hold for both constant and increasing BSs.
[…]
where $Q_{1},Q_{2},\tilde{Q}_{1}$, and $\tilde{Q}_{2}$ are constants that do not depend on $T$.
@@ 3.3 Case (iii): Increasing BS; Warm-up Decaying LR
In this case, we consider a BS $(b_{t})_{t}$ and an LR $(\eta_{t})_{t}$ such that $b_{t}\leq b_{t+1}$ and $\eta_{t}\leq\eta_{t+1}$ for $(t\leq T_{w}-1)$ and $\eta_{t+1}\leq\eta_{t}$ for $(t\geq T_{w})$. As examples of an increasing warm-up LR, we consider an exponential growth LR and a polynomial growth LR, both increasing every $K^{\prime}$ steps. We set $K$, as defined in Case (ii), to be $lK^{\prime}$, where $l\in\mathbb{N}$ (thus $K>K^{\prime}$). Namely, we consider a setting in which the BS is increased every $l$ times the LR is increased. To formulate examples of an increasing LR, we define $M^{\prime}\coloneq\lfloor\frac{T}{K^{\prime}}\rfloor$ and formalize the LR: for every $m\in\{0,\ldots,M^{\prime}-1\},t\in S^{\prime}_{m}\coloneq[mK^{\prime},(m+1)K^{\prime})\cap\mathbb{Z}$, $\text{Exponential Growth LR:}\ \eta_{t}\coloneq\eta_{0}\delta^{m\left\lceil\frac{t}{(m+1)K^{\prime}}\right\rceil},$ (7) $\text{Polynomial Growth LR:}\ \eta_{t}\coloneq\left(sm\left\lceil\frac{t}{(m+1)K^{\prime}}\right\rceil+\eta_{0}\right)^{q},$ (8) where $s>0$ and $q>1$. Furthermore, we choose $\gamma,\delta>1$, and $l\in\mathbb{N}$ such that $\delta^{2l}<\gamma$ holds. Additionally, we set $l_{w}\in\mathbb{N}$ such that $T\geq T_{w}\coloneq l_{w}K^{\prime}\geq lK^{\prime}$. The examples of an increasing BS used in this case are an exponential growth BS (5) and a polynomial growth BS (6). As examples of a warm-up LR, we use LRs that are increased using the exponential growth LR (7) and the polynomial growth LR (8) corresponding respectively to the exponential and polynomial growth BSs for the first $T_{w}$ steps and then decreased using the constant LR (1), the diminishing LR (2), the cosine annealing LR (3), or the polynomial decay LR (4) for the remaining $T-T_{w}$ steps. Note that $\eta_{\max}\coloneq\eta_{T_{w}-1}$. A more detailed version of Theorem 3.4 is provided in Appendix D.4.
Theorem 3.4.
We consider BSs (5) and (6) together with warm-up LRs (7) and (8) with decay parts given by (1), (2), (3), or (4) under the assumptions stated in Lemma 3.1. Then, the following results hold for both constant and increasing BSs: (I) Decay part [Diminishing (2)]: $\|G_{T}\|^{2}=O((\sqrt{T+1}-\sqrt{T_{w}+1})^{-1})$; (II) Decay part [Otherwise (1), (3), (4)]: $\|G_{T}\|^{2}=O((T-T_{w})^{-1})$.
@@ 3.4 Case (iv): Constant BS; Warm-up Decaying LR
We consider a BS $(b_{t})_{t}$ and an LR $(\eta_{t})_{t}$ such that $b_{t}=b$ and $\eta_{t}\leq\eta_{t+1}$ for $(t\leq T_{w}-1)$ and $\eta_{t+1}\leq\eta_{t}$ for $(t\geq T_{w})$. As examples of a warm-up LR, we use the exponential growth LR (7) and the polynomial growth LR (8) for the first $T_{w}$ steps and then the constant LR (1), the diminishing LR (2), the cosine annealing LR (3), or the polynomial decay LR (4) for the remaining $T-T_{w}$ steps. The other conditions are the same as those in Case (iii). A more detailed version of Theorem 3.5 is provided in Appendix D.5.
Theorem 3.5.
We consider a constant BS $b_{t}=b>0$ and warm-up LRs (7) and (8) with decay parts given by (1), (2), (3), or (4) under the assumptions stated in Lemma 3.1. Then, we obtain the following results: (I) Decay part [Diminishing (2)]: $\|G_{T}\|^{2}=O(\log\frac{T}{T_{w}}(\sqrt{T+1}-\sqrt{T_{w}+1})^{-1})$, (II) Decay part [Otherwise (1), (3), (4)]: $\|G_{T}\|^{2}=O((T-T_{w})^{-1}+b^{-1})$.
@@ 4 Numerical Experiment
We experimentally evaluated the performance of RSGD for the two types of BSs and various types of LRs introduced in Section 3. The experiments were run on an iMac (Intel Core i5, 2017) running the macOS Ventura operating system (ver. 13.7.1). The algorithms were written in Python (3.12.7) using the NumPy (1.26.0) and Matplotlib (3.9.1) packages. The Python code is available at https://github.com/iiduka-researches/RSGD_acml2025.git. We set $p=2.0$ in (4) and $\eta_{\min}:=0$. In Cases (i) and (ii), we used an initial LR $\eta_{\max}$ selected from $\{0.5,0.1,0.05,0.01,0.005\}$. In Case (ii), we set $K=1000,\gamma=3.0$, and $a=c=2.0$. All the plots of the objective function values presented in this section are provided in Appendices A.2 and E. Those for Cases (iii) and (iv) are provided in Appendix G.
@@ 4.1 Principal Component Analysis
Figure 1: Norm of the gradient of the objective function versus the number of iterations for LRs (1), (2), (3), and (4) on COIL100 dataset (PCA).
Figure 2: Norm of the gradient of the objective function versus the number of iterations for LRs (1), (2), (3), and (4) on MNIST dataset (PCA).
We can formulate the PCA problem as an optimization problem on the Stiefel manifold (Kasai et al., 2019); for a given dataset $\{x_{j}\}_{j=1,\ldots,N}\subset\mathbb{R}^{n}$ and $r$ $(\leq n)$, $\text{minimize}_{U\in\mathrm{St}(r,n)}f(U)\coloneq\frac{1}{N}\sum_{j=1}^{N}\|x_{j}-UU^{\top}x_{j}\|^{2}$, where $\mathrm{St}(r,n)\coloneq\{U\in\mathbb{R}^{r\times n}\mid U^{\top}U=I_{n}\}$. We set $r=10$ and used the COIL100 (Nene et al., 1996) and MNIST (LeCun et al., 1998) datasets. The Columbia Object Image Library (COIL100) dataset contains $7200$ color camera images of $100$ objects ($72$ poses per object) taken from different angles. We resized the images to $32\times 32$ pixels and transformed each one into a $1024$ $(=32^{2})$ dimensional vector. Hence, we set $(N,n,r)=(7200,1024,10)$. The MNIST dataset contains $60,000$ $28\times 28$-pixel grayscale images of handwritten digits 0 to $9$. We transformed each image into a $784$ $(=28^{2})$ dimensional vector and normalized each pixel to the range $[0,1]$. Hence, we set $(N,n,r)=(60000,784,10)$. Furthermore, we used a constant BS with $b_{t}\coloneq 2^{10}$, an exponential growth BS with an initial value $b_{0}\coloneq 3^{5}$, and a polynomial growth BS with an initial value $b_{0}\coloneq 30$.
@@ 4.2 Low-rank Matrix Completion
Figure 3: Norm of the gradient of the objective function versus the number of iterations for LRs (1), (2), (3), and (4) on MovieLens-1M dataset (LRMC).
@@
Figure 4: Norm of the gradient of the objective function versus the number of iterations for LRs (1), (2), (3), and (4) on Jester dataset (LRMC).
@@
The LRMC problem involves completing an incomplete matrix $Z=(z_{1},\ldots,z_{N})\in\mathbb{R}^{n\times N}$; $\Omega$ denotes a set of indices for which we know the entries in $Z$. For $a\in\mathbb{R}^{n}$, we define $P_{\Omega_{i}}(a)$ such that the $j$-th element is $a_{j}$ if $(i,j)\in\Omega$ and 0 otherwise. For $U\in\mathbb{R}^{n\times r},z\in\mathbb{R}^{n}$, $q_{j}(U,z):=\mathrm{argmin}_{a\in\mathbb{R}^{r}}\|P_{\Omega_{j}}(Ua-z)\|$. We can now formulate the LRMC problem as the following optimization problem on the Grassmann manifold (Boumal and Absil, 2015): $\text{minimize}_{U\in\mathrm{Gr}(r,n)}f(U)\coloneq\frac{1}{N}\sum_{j=1}^{N}\|P_{\Omega_{j}}(Uq_{j}(U,z_{j})-z_{j})\|^{2}$, where $\mathrm{Gr}(r,n)\coloneq\text{St}(r,n)/\text{O}(r)$. We set $r=10$ and used the MovieLens-1M (Harper and Konstan, 2016) and Jester datasets (Goldberg et al., 2001). The MovieLens-1M dataset contains $1,000,209$ ratings given by $6040$ users on $3952$ movies. Every rating lies in $[0,5]$. We normalized each rating to the range $[0,1]$. Hence, we set $(N,n,r)=(3952,6040,10)$. The Jester dataset contains ratings of $100$ jokes from $24,983$ users. Every rating is bounded by the range [10,10][-10,10]. Hence, we set $(N,n,r)=(24983,100,10)$. Furthermore, we used a constant BS $b_{t}\coloneq 2^{8}$, an exponential growth BS with an initial value of $b_{0}\coloneq 3^{4}$, and a polynomial growth BS with an initial value of $b_{0}\coloneq 14$.
@@
The performances in terms of the gradient norm of the objective function versus the number of iterations for LRs (1), (2), (3), and (4) on the COIL100, MNIST, MovieLens-1M, and Jester datasets are plotted in Figures 1, 2, 3, and 4, respectively. Achieving a small gradient norm was better with an increasing BS than with a constant BS. Among the increasing BSs, the exponential growth BS outperformed the polynomial BS. Because the magnitude of increase in the exponential BS is $O(\gamma^{m})$ and that in the polynomial BS is $O(m)$, these numerical results indicate that a larger rate of increase in BS leads to better performance.
@@
4.3 Comparison of Computational Time between Constant BS and Increasing BS versus SFO Complexity